OpenStudy (anonymous):

Simplify square root parenthesis 1 minus cosine theta parenthesis times parenthesis 1 plus cosine theta parenthesis divided by cosine squared theta. square root sine theta ±sin Θ<------ MY ANSWER ±cos Θ |tan Θ| CAN ANYONE PLEASE CHECK IF I GOT THIS RIGHT?!!!

OpenStudy (anonymous):

@e.mccormick

OpenStudy (e.mccormick):

\(\sqrt{1 - \cos\theta} \cdot \dfrac{1 +\cos \theta }{ \cos^2 \theta}\) That?

OpenStudy (e.mccormick):

Or looking at where you said parenthesis it could be... \(\dfrac{\sqrt{(1 - \cos\theta)(1 +\cos \theta )} }{ \cos^2 \theta}\)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ganeshie8 @ganeshie8

OpenStudy (e.mccormick):

Still need to know what the equation is supposed to be... You just wrote parenthesis so I di not know what opens and closes.